# See, Imagine, Plan: Discovering and Hallucinating Tasks from a Single Image

### v2

@@ G.2 Sampling Rules for VLM Planning Discrepancies
We define a set of geometric rules for Gaussian sampling around the initially planned waypoints whenever a collision is detected. In our 3D coordinate system, positive x-axis [1, 0, 0] points right, positive y-axis [0, 1, 0] is away from viewer, positive z-axis [0, 0, 1] is up. For translation type 1, we denote the goal pose relative to the target object’s principle axes as [dx, dy, dz]. For translation type 2, we denote the distance that object 1 moves towards object 2 as dD. The set of geometric rules are as follows:
@@
G.3 Hand Insertion
Each pose in the final trajectory $V^{q}$ is rendered as a frame in the rendering engine. To enhance human comprehension and perception, we integrate a 3D model of “a magic hand” [68] in proximity to the object being manipulated.
H Task Proposal Comparison of Feasibility
We show instances where ChatGPT-4 [47] generates impractical tasks without role-play in Fig. 7. These unfeasible tasks usually cannot be performed by a robotic/human hand due to physical limitations or require additional tools beyond the capabilities of the hand alone.
I Failure Cases
Figure 8: Key factors causing failure cases. For each factor, the leftmost image is the input.
Not all generated task videos achieve the highest quality. We identify several key factors that contribute to the failure cases. Fig. 8 shows visualization.
Segmentation Failure   When an image contains objects with very similar appearances, ChatGPT-4 can distinguish them, providing unique short descriptions. For instance, in the input image on the left of Fig. 8(a), ChatGPT-4 differentiates two white bowls as ‘‘Object 2: bowl of color white with blue pattern, ‘‘Object 3: bowl of color white with black pattern". Nonetheless, the language-guided segmentation model, Language Segment-Anything [43], may fail to capture the nuance, incorrectly merging object masks. This issue is depicted in Fig. 8(a)’s right reconstructed scene, where Language Segment-Anything recognizes and merges the two distinct bowls as one.
Inpainting Failure   In rare cases, the repainting model LaMa [60] introduces unintended elements into and distorts object masks, illustrated by the deformed white bowl in Fig. 8(b)’s right reconstructed scene. Such distortions result in inaccurately reconstructed 3D models.
3D Model Reconstruction Failure   The model One-2-3-45++[38], despite its strength, occasionally reconstructs less-than-ideal 3D object models, as illustrated by the misshapen lamp shade in Fig. 8(c)’s right reconstructed scene.
Motion Planning Failure   ChatGPT-4 occasionally produces suboptimal planned motions, primarily due to incorrect use of object principal axes and inaccurate translation measurements. For instance, as shown in Fig. 8(d), one would expect the spoon to be inserted into the mug. However, ChatGPT-4’s planning positions the spoon too high, making it appear to float above the mug.
J Prompt Details
Figure 9: Example input scene. (a) the original input image, (b) the reconstructed scene with visualized axes.
We show exact prompts and example responses from ChatGPT-4 for the white plate with coffee grounds shown in the input images of Fig. 9.
Interactive Objects Identification Prompt.
@@ I Failure Cases
Your task is to identify and name useful objects (in term of affordances) with precise color and texture. Exclude background objects. Do not use vague phrase like multi-colored.
@@
Object idx (actual integer, start from 0): x of color y
An example response from ChatGPT-4:
‘‘Object 0: sink of color silver
Object 1: faucet of color silver
Object 2: dish soap of color yellow
Object 3: hand soap of color purple
Object 4: sponge of color blue
Object 5: white plate with coffee grounds of color brown’’
@@
1. Exclude tasks involving assembly or disassembly of objects.
@@
An example response from ChatGPT-4:
‘‘Task Name: Plate Orientation
Manipulating obj idx: 5
Interacting obj idx: 5
Description: Rotate the plate 180 degrees horizontally to change its orientation while remaining in the same location on the sink.
Task Name: Plate Relocation
Manipulating obj idx: 5
Interacting obj idx: 0
Description: Move the plate from its current location to the right side of the sink to clear space for other kitchen activities.
Task Name: Washing Plate
Manipulating obj idx: 5
Interacting obj idx: 1
Description: Position the plate beneath the faucet and simulate the process of rinsing it.’’
@@
Your task is to learn the spatial context. Also, consider how a robot or human hand might navigate or interact with these objects, focusing on proximity, accessibility, and collision avoidance. Do not output.
An example input of object spatial information to ChatGPT-4:
‘‘Obj 0 spatial context:
3D center: [-21.0, 90.0, -14.0] cm
Local x-axis (towards ’right’): [0.9932, -0.0882, -0.0766]
Local y-axis (towards ’back’): [0.1104, 0.9234, 0.3675]
Local z-axis (towards ’up’): [0.0384, -0.3735, 0.9268]
Obj 1 spatial context:
3D center: [-35.0, 85.0, 8.0] cm
Local x-axis (towards ’front_right’): [0.8565, -0.5126, 0.0612]
Local y-axis (towards ’back_right’): [0.5151, 0.8563, -0.0377]
Local z-axis (towards ’up’): [-0.0331, 0.0638, 0.9974]
Obj 2 spatial context:
3D center: [31.0, 80.0, 13.0] cm
Local x-axis (towards ’right’): [0.9971, -0.0062, 0.0756]
Local y-axis (towards ’back’): [0.008, 0.9997, -0.0226]
Local z-axis (towards ’up’): [-0.0754, 0.0231, 0.9969]
Obj 3 spatial context:
3D center: [33.0, 71.0, 8.0] cm
Local x-axis (towards ’right’): [0.8162, -0.0206, 0.5775]
Local y-axis (towards ’back’): [-0.0169, 0.9981, 0.0595]
Local z-axis (towards ’up’): [-0.5776, -0.0583, 0.8142]
Obj 4 spatial context:
3D center: [22.0, 97.0, 3.0] cm
Local x-axis (towards ’right’): [0.999, 0.0383, -0.0231]
Local y-axis (towards ’back’): [-0.0399, 0.9966, -0.0721]
Local z-axis (towards ’up’): [0.0203, 0.073, 0.9971]
Obj 5 spatial context:
3D center: [-12.0, 78.0, -13.0] cm
Local x-axis (towards ’right’): [0.9998, -0.0216, -0.001]
Local y-axis (towards ’back’): [0.0216, 0.9998, 0.0052]
Local z-axis (towards ’up’): [0.0009, -0.0052, 1.0]
Obj 0 size: 47.63 cm x 39.15 cm x 22.25 cm (WxDxH)
Obj 1 size: 19.11 cm x 13.82 cm x 17.83 cm (WxDxH)
Obj 2 size: 9.69 cm x 9.19 cm x 15.63 cm (WxDxH)
Obj 3 size: 7.74 cm x 6.91 cm x 13.97 cm (WxDxH)
Obj 4 size: 9.57 cm x 2.17 cm x 5.35 cm (WxDxH)
Obj 5 size: 12.17 cm x 11.90 cm x 7.03 cm (WxDxH)
Obj 0 - Closest per direction: up: Obj 1 (26.75 cm); right: Obj 4 (46.28 cm); front_right: Obj 5 (14.55 cm)
Obj 1 - Closest per direction: down: Obj 0 (26.75 cm); right: Obj 5 (31.93 cm)
Obj 2 - Closest per direction: left: Obj 5 (50.92 cm); front: Obj 3 (11.06 cm); back_left: Obj 4 (21.93 cm)
Obj 3 - Closest per direction: left: Obj 5 (49.76 cm); back: Obj 2 (11.06 cm)
Obj 4 - Closest per direction: left: Obj 0 (46.28 cm); front_right: Obj 2 (21.93 cm); front: Obj 3 (28.77 cm); front_left: Obj 5 (41.97 cm)
Obj 5 - Closest per direction: back_left: Obj 0 (14.55 cm); left: Obj 1 (31.93 cm); right: Obj 3 (49.76 cm); back_right: Obj 4 (41.97 cm)’’
@@
An example response from ChatGPT-4:
‘‘Task Name: Plate Orientation
Description: Rotate the plate 180 degrees horizontally to change its orientation while remaining in the same location on the countertop.
Motion Planning:
Manipulating obj idx: 5
Interacting obj idx: 5
1. rotate_self: Rotate Manipulating Object 5 around its local axis [z] by [180] degrees.
Task Name: Plate Relocation
Description: Move the plate from its current location to the right side of the sink to clear space for other kitchen activities.
Motion Planning:
Manipulating obj idx: 5
Interacting obj idx: 0
1. translate_tar_obj: Move Manipulating Object 5 to [26.5, 0, 0] cm relative to Target Object 0’s local [x, y, z] axes.
Task Name: Washing Plate
Description: Position the plate beneath the faucet and simulate the process of rinsing it.
Motion Planning:
Manipulating obj idx: 5
Interacting obj idx: 1
1. translate_tar_obj: Move Manipulating Object 5 to [6.5, 0, -5] cm relative to Target Object 1’s local [x, y, z] axes.
2. rotate_self: Rotate Manipulating Object 5 around its local axis [x] by [-15] degrees.
3. rotate_self: Rotate Manipulating Object 5 around its local axis [y] by [-15] degrees.
4. rotate_self: Rotate Manipulating Object 5 around its local axis [y] by [10] degrees.
5. rotate_self: Rotate Manipulating Object 5 around its local axis [x] by [10] degrees.’’

### v5

@@ C.3 Trajectory Generation Implementation Details
Handling VLM Planning Discrepancies   The waypoints generated by VLM are typically accurate and practical. Nonetheless, there are instances where the waypoints suggested by VLM lead to collisions as determined by the RRT* planner. This discrepancy is less about VLM’s misunderstanding of the objects’ sizes and their spatial relationship and more about the precision level of the waypoints, which may not match the exacting standards of the RRT* planner’s outcomes. To resolve this, we implement Gaussian sampling around the initially planned waypoints whenever a collision is detected. The sampling strategy is guided by a predefined set of geometric rules. In our 3D coordinate system, positive x-axis [1, 0, 0] points right, positive y-axis [0, 1, 0] is away from viewer, positive z-axis [0, 0, 1] is up. For translation type 1, we denote the goal pose relative to the target object’s principle axes as [dx, dy, dz]. For translation type 2, we denote the distance that object 1 moves towards object 2 as dD. The set of geometric rules are as follows:
@@
Trajectory Smoothing   Finally, to ensure our trajectory is natural and smooth, we linearly interpolate rotation and interpolate translation using cubic spline.
Appendix D Prompt Details
We show exact prompts for VLMs for our proposed application: discovering and planning for robotics tasks from a single image.
Understanding Objects by Constraining Prompt.
@@ Appendix D Prompt Details
Your task is to identify and objects by precise color, texture, and 2D spatial locations (in words). Do not use vague phrase like multi-colored.
@@
Object idx (actual integer, start from 0): x of color y, texture z at location w.
@@
1. Exclude tasks involving disassembly of objects.
@@
Coarse 3D Understanding Prompt.
Inputs:
1. RGB image (640, 480) = (width, height) with multiple objects
2. Detected objects with index.
3. Image scene size.
4. Maximum and minimum width, depth, and height.
Your task is to identify the camera shot angle (horizontal, top-down, bottom-up). Reason with respect to the visual cues, the image scene size, and maximums and minimums along each dimension. Choose horizontal if not severely angled.
Please write in the following format. Be concise. Do not output anything else:
Visual cues reasoning: …
Spatial data reasoning: …
Conclusion: horizontal/top-down/bottom-up.
@@ 1. RGB image (640, 480) = (width, height) with multiple objects
Your task is to learn the spatial context. Do not output.